Assume that the total cost of a competitive firm is given by,TC=6Q2-16Q+12, and the total revenue of the firm is given by, TR=2Q2

Then find

A.The firms profit maximization level of output.

B.Profit

C.does the firm generate profit or incur loss at 5 unit of output?


1
Expert's answer
2022-03-23T14:01:37-0400

"TC=6Q^2-16Q+12"

MC=12Q-16

"TR=2Q^2"

"P= \frac{TR}{Q}= \frac{2Q^2}{Q}=2Q"

For profit maximization, MC= P

2Q= 12Q-16

Q= 1.6

P= 2(1.6)=3.2

"\Pi= TR- TC"

=2("1.6^2)" - ["6(1.6^2)-16(1.6)+12]"

5.12-1.76= 3.36

c) At Q=5

"\Pi=" "2(5^2) -[6(5^2)-16(5)+12]"

= 58

It makes a profit of 58